find an equation of the line containing the given pair of points

(-5,-1) and ( -2,-8)

Answer by unlockmath(1122)

Hello,
First, let's find the slope of the line using the slope formula (y-y)/(x-x)
(-5,-1) and ( -2,-8) It will look like this:
-1-(-8)/-5-(-2) will give us:
7/-3 or -7/3 Now we can use the point-slope formula which is:
y-y=m(x-x) Plug in either of the points and we have: (Note: m is slope)
y-(-1)=-7/3(x-(-5) Simpilfy this to:
y+1= -7/3x - 35/3 Subtract 1 both sides which gives us:
y=-7/3(x) - 38/3 This can be rewritten as:
y = -2 1/3(x) - 12 2/3
RJ Toftness
www.math-unlock.com